ADNAN MUHAMMED

Grade 12,

two balls of equal mass undergo head on collision while each was moving with speed 6m/s. if the coefficient of restitution is 1/3, the speed of each ball after impact will be ??? explain it to me its answer is 2 m/s...

two balls of equal mass undergo head on collision while each was moving with speed 6m/s. if the coefficient of restitution is 1/3, the speed of each ball after impact will be ??? explain it to me its answer is 2 m/s...

Grade:11

3 Answers

DR STRANGE
7 years ago

Hi VAIBHAV,

USE THIS CONCEPT,

relative velocity of separartion = e*relative velocity of approach 

 

AS THE BALLS ARE EQUAL AND ARE UNDERGOING HEAD ON COLLISION

(while each was moving with equal speed)

The speed of the balls will be same after collision

relative velocity of approach is 6-(-6)=12m/s., coefficient of restitution is 1/3

put this in the above equation,

you will get relative velocity of separation=4m/s

as the balls will move in opposite direction after collision, the velocity of one ball will be 2m/s and -2m/s for other. (i.e relative velocity=2-(-2)=4m/s)

hence speed of each ball will be 2m/s (considering inly magnitude of the velocities)

Plz approve if you are satisfied :)

Vaibhav
7 years ago
i want to ask u dr strange that cant we simply multiply the velocity with the coefficient of restitution that is 1/3 x 6 =2. ... (plz approve if it is possible) xD
DR STRANGE
7 years ago
Approved Vaibhav sir :P,
yes, this is possible, but conditions:
1.both ball should be same
2.it shoulld be head on collision
3. there initial speed should be same.